### v1

@@ Abstract
Sequential recommendation is one of the important branches of recommender system, aiming to achieve personalized recommended items for the future through the analysis and prediction of users’ ordered historical interactive behaviors. However, along with the growth of the user volume and the increasingly rich behavioral information, how to understand and disentangle the user’s interactive multi-intention effectively also poses challenges to behavior prediction and sequential recommendation. In light of these challenges, we propose a Contrastive Learning sequential recommendation method based on Multi-Intention Disentanglement (MIDCL). In our work, intentions are recognized as dynamic and diverse, and user behaviors are often driven by current multi-intentions, which means that the model needs to not only mine the most relevant implicit intention for each user, but also impair the influence from irrelevant intentions. Therefore, we choose Variational Auto-Encoder (VAE) to realize the disentanglement of users’ multi-intentions, and propose two types of contrastive learning paradigms for finding the most relevant user’s interactive intention, and maximizing the mutual information of positive sample pairs, respectively. Experimental results show that MIDCL not only has significant superiority over most existing baseline methods, but also brings a more interpretable case to the research about intention-based prediction and recommendation.
@@ 1 Introduction
Nowadays, recommender system has gradually become an indispensable part of people’s lives, and the more realistic and logical sequence recommendation has become the key to provide personalized needs. As a popular research direction, sequential recommendation is not only for product recommendation, but also can be extended to other fields, such as intention understanding and prediction. The main reason is that the definition of sequential recommendation is not absolute, such as the task of recommending the next product to the user based on the browsing and purchasing records, but we can also consider it as a prediction task based on a series of consecutive historical interactive sequences of the user, and analyze the potential intention of users to generate the most probable interactions that will occur at the next moment, so that the method can be applied to real-world datasets in different scenarios.
The difference between prediction and recommendation is that recommendation usually does not recommend the interacted items to the user again, but the behavior sequence may contain a large number of repeated interactions, which leads to a wider scope of prediction than recommendation. Besides, due to the real-time progression of interactive intention, we believe that the closer the interactive behavior is to the current moment, the more it can influence the prediction of the next moment. Fig.1 illustrates the sequence of four users’ interactive behaviors, and it can be seen that there are a variety of coherent behaviors that contain certain fixed-pairing items. In particular, the most recent interacted item has the highest weight in the prediction task, but this is not absolute. As shown in this figure, it can be found that the first and last users have different interacted tail items, but the items to be predicted are the same. Therefore, it is reasonable to assume that although the two users have different behaviors at this point in time, their intentions are indeed close to each other, which reflects the importance of intention understanding and disentanglement.
@@
The core task of sequence recommendation is to predict whether the user will interact with the candidate item at the next moment, and different scenarios will be handled in different ways. For example, in the graph-based recommendation scenario, we can regard it as a link prediction problem. Through the current research on sequential recommendation, it can be seen that most of the work focuses on the feature construction of the user and the item, but instead does not consider the historical behavior that contains rich potential information, and ignores the full mining of the user’s interactive intention. Meanwhile, the input features of most models in Computer Vision (CV) and Natural Language Processing (NLP) are dense features, while the inputs of recommendation models are mostly sparse features such as id information, so how to design data enhancement methods for sparse features is the key to optimize the sequential recommendation task.
@@ 2.1 Sequential Recommendation
Unlike traditional recommendation based on collaborative filtering of users or items, sequential recommendation pays more attention to the interaction sequence of users’ behavior history, focusing on mining users’ real-time interests and interactive intentions. The earliest sequential recommendation started with statistical machine learning methods, based on Markov chain [1] to predict the next clicked item through the user’s historical sequence; combining the first-order Markov chain and matrix factorization [2] not only focuses on the user’s history, but also simulates the user’s potential preferences; combining similarity computation in the Markov chain [3] to a certain extent compensates for the shortcomings of the first two models. To a certain extent, at that time, the first two models are more effective and more efficient, which achieved better results.
With the rise of deep learning, researchers began to apply it to sequential recommendation because of its powerful ability to handle massive data and complex problems. For example, GRU4Rec [4], which improves RNN, alleviates the former problem of gradient vanishing or gradient explosion; Caser [5], based on CNN, enhances feature associations between items by splicing the convolutional results in different directions; SASRec [6] based on the self-attention mechanism [7], fully exploits the user’s interest weight in the current interaction; and BERT4Rec [8] consists of the bidirectional Transformer-based structure, which demonstrates its powerful processing and computing capabilities in the face of large-scale data.
However, most of the existing algorithms for sequence recommendation lack the analysis of the user’s dynamic interests and the interpretation of interaction intentions. As a result, we build on the above and conduct a series of work that incorporates the intent during interaction behavior.
@@ 3.3 Contrastive Learning
$Score\bigl{(}f(x),f(x^{+})\bigr{)}\gg Score\bigl{(}f(x),f(x^{-})\bigr{)}.$ (4)
@@ 5.1.3 Baseline Methods
  • GRU4Rec [4]: Modeling progressive dependencies of user sequence elements via GRUs.
@@ 5.5 Intention Visualization (RQ4)
Fig.LABEL:fig6 illustrates the eight stages of a certain training process in which the sequence and intention features of users are clustered after dimensionality reduction to the two-dimensional plane, where the large dark blue dots indicate the disentangled $k$-intents, and this time $k=8$ is used as an example. The remaining small dots of eight different colors indicate the sequence features of the users in one batch-size, the same color represents the same clusters, and the features of the users with the same clusters means they have a higher degree of similarity.
As shown in the figure, the beginning of the training of all the user’s sequence features present a number of clusters of different sizes and randomly distributed; subsequently, the user features began to gradually spread, as the gradient of the loss decreases, each user began to draw closer to its similar intentions; at the end, when the iteration tends to stabilize, all users basically have found the closest to their own intention representation, and constitute the $k$ clusters. At this point, it is reasonable to assume that users within the same cluster have the most similar intention, and they are more likely to produce similar even same interaction behaviors in the next moment.
@@ 6 Conclusion
In this paper, we propose a contrastive learning method MIDCL based on multi-intention disentanglement for sequential recommendation, which enables the prediction of the next item in a sequence by mining user’s intention. First, the behavioral sequences are embedded by Transformer encoder; second, multi-intentions of users are disentangled based on VAE in the potential space with a parameter-sharing Transformer encoder, whose dimensions are used as the number of user’s latent intentions; finally, we propose two contrastive learning paradigms to optimize the feature representations of the intentions and sequences, respectively, to weaken the negative impact of irrelevant intentions. The experimental results on four real-world datasets show that MIDCL achieves the optimal results in all baseline methods and validates the contribution of the intention disentanglement and contrastive learning modules, and we also conduct the visualization of the intentions as a way to demonstrate the feasibility and interpretability of intention disentanglement module.

### v2

@@ Abstract
Sequential recommendation is one of the important branches of recommender system, aiming to achieve personalized recommended items for the future through the analysis and prediction of users’ ordered historical interactive behaviors. However, along with the growth of the user volume and the increasingly rich behavioral information, how to understand and disentangle the user’s interactive multi-intention effectively also poses challenges to behavior prediction and sequential recommendation. In light of these challenges, we propose a Contrastive Learning sequential recommendation method based on Multi-Intention Disentanglement (MIDCL). In our work, intentions are recognized as dynamic and diverse, and user behaviors are often driven by current multi-intentions, which means that the model needs to not only mine the most relevant implicit intention for each user, but also impair the influence from irrelevant intentions. Therefore, we choose Variational Auto-Encoder (VAE) to realize the disentanglement of users’ multi-intentions. We propose two types of contrastive learning paradigms for finding the most relevant user’s interactive intention, and maximizing the mutual information of positive sample pairs, respectively. Experimental results show that MIDCL not only has significant superiority over most existing baseline methods, but also brings a more interpretable case to the research about intention-based prediction and recommendation.
@@ 1 Introduction
Nowadays, recommender system has gradually become an indispensable part of people’s lives, meanwhile, the more realistic and logical sequential recommendation are becoming the key to provide personalized needs. As a popular research direction, sequential recommendation is not only for product recommendation, but also can be extended to other fields, such as intention understanding and prediction. The main reason is that the definition of sequential recommendation is not absolute, such as the task of recommending the next product to the user based on the browsing and purchasing records. But we can also consider it as a prediction task based on a series of consecutive historical interactive sequences of the user, and analyze the potential intention of users to generate the most probable interactions that will occur at the next moment, so that the method can be applied to real-world datasets in different scenarios.
The difference between prediction and recommendation is that recommendation usually does not recommend the interacted items to the user again, but the behavior sequence may contain a large number of repeated interactions, which leads to a wider scope of prediction than recommendation. Due to the real-time progression of interactive intention, we believe that the closer the interactive behavior is to the current moment, the more it can influence the prediction of the next moment. Fig.1 illustrates the sequence of four users’ interactive behaviors, and it can be seen that there are a variety of coherent behaviors that contain certain fixed-pairing items. In particular, the most recent interacted item has the highest weight in the prediction task, but this is not absolute. As shown in this figure, it can be found that the first and last users have different interacted tail items, but the items to be predicted are the same. Therefore, it is reasonable to assume that although the two users have different behaviors at this point in time, their intentions are indeed close to each other, which reflects the importance of intention understanding and disentanglement.
@@
The core task of sequential recommendation is to predict whether the user will interact with the candidate item at the next moment, and different scenarios will be handled in different ways. For example, in the graph-based recommendation scenario, we can regard it as a link prediction problem. Through the current research on sequential recommendation, it can be seen that most of the work focuses on the feature construction of the user and the item, but instead does not consider the historical behavior that contains rich potential information, and ignores the full mining of the user’s interactive intention. Meanwhile, the input features of most models in Computer Vision (CV) and Natural Language Processing (NLP) are dense features, while the inputs of recommendation models are mostly sparse features such as id information, so how to design data enhancement methods for sparse features is the key to optimize the sequential recommendation task.
@@ 2.1 Sequential Recommendation
Unlike traditional recommendation based on collaborative filtering of users or items, sequential recommendation pays more attention to the interaction sequence of users’ behavior history, focusing on mining users’ real-time interests and interactive intentions. The earliest sequential recommendation started with statistical machine learning methods, based on Markov chain [1] to predict the next clicked item through the user’s historical sequence; combining the first-order Markov chain and matrix factorization [2] not only focuses on the user’s history, but also simulates the user’s potential preferences; combining similarity computation in the Markov chain [3] to a certain extent compensates for the shortcomings of the first two models.
With the rise of deep learning, researchers began to apply it to sequential recommendation because of its powerful ability to handle massive data and complex problems. For example, GRU4Rec [4], which improves Recurrent Neural Network (RNN), alleviates the former problem of gradient vanishing or gradient explosion; Caser [5], based on Convolution Neural Network (CNN), enhances feature associations between items by splicing the convolutional results in different directions; SASRec [6] based on the self-attention mechanism [7], fully exploits the user’s interest weight in the current interaction, and TiSASRec [8] incorporates consideration of the time intervals between interactions based on  [6]. BERT4Rec [9] consists of the bidirectional Transformer-based structure, which demonstrates its powerful processing and computing capabilities in the face of large-scale data.
Nevertheless, most of the existing algorithms for sequential recommendation lack the analysis of the user’s dynamic interests and the interpretation of interaction intentions. In some studies, Graph Neural Network (GNN) has been employed to deal with heterogeneous networks [10, 11, 12, 13, 14], and some others try to discover latent user interests through contextual information  [15, 16, 17] or interaction time-aware analysis  [18, 19, 20], but we hope to mine the potential multiple intentions of users from the interaction sequences themselves. As a result, we build on the above and conduct a series of work that incorporates the intention during interaction behavior.
@@ 3.3 Contrastive Learning
$Score\bigl{(}f(x),f(x^{+})\bigr{)}\gg Score\bigl{(}f(x),f(x^{-})\bigr{)},$ (4)
@@ 5.1.3 Baseline Methods
  • GRU4Rec [4]: Modeling progressive dependencies of user sequence elements via Gated Recurrent Unit (GRU)s.
@@ 5.5 Intention Visualization (RQ4)
Fig.LABEL:fig6 illustrates the different stages of a certain training process in which the sequence and intention features of users are clustered after dimensionality reduction to the two-dimensional plane, where the large dark blue dots indicate the disentangled $k$-intents, and this time $k=8$ is used as an example. The remaining small dots of eight different colors indicate the sequence features of the users in one batch-size, the same color represents the same clusters, and the features of the users with the same clusters means they have a higher degree of similarity.
As shown in the figure, (a) represents the beginning of the training of all the user’s sequence features present a number of clusters of different sizes and randomly distributed; subsequently, from (b) to (e), the user features began to gradually spread; as the gradient of the loss decreases, each user began to draw closer to its similar intentions from (f) to (h); at the end in (i), when the iteration tends to stabilize, all users basically have found the closest to their own intention representation, and constitute the $k$ clusters. With given data, it is reasonable to assume that users within the same cluster have the most similar intention, and they are more likely to produce similar even same interaction behaviors in the next moment.
@@ 6 Conclusion
In this paper, we propose a contrastive learning method MIDCL based on multi-intention disentanglement for sequential recommendation, which enables the prediction of the next item in a sequence by mining user’s intention. First, the behavioral sequences are embedded by Transformer encoder; second, multi-intentions of users are disentangled based on VAE in the potential space with a parameter-sharing Transformer encoder, whose dimensions are used as the number of user’s latent intentions; finally, we propose two contrastive learning paradigms to optimize the feature representations of the intentions and sequences, respectively, to weaken the negative impact of irrelevant intentions. The experimental results on four real-world datasets show that MIDCL achieves the optimal results in all baseline methods and validates the contribution of the intention disentanglement and contrastive learning modules. To explore changes in intention with training, we also conduct the visualization of the intentions as a way to demonstrate the feasibility and interpretability of intention disentanglement module.
@@
[…]
Acknowledgements
This paper is supported by the National Key Research and Development Program of China (314) and the Independent Project of National Key Laboratory (2024-SKL-005).